OpenStudy (anonymous):

how can you find the roots of this polynomial equation:2x^4-7x^#+21x^2+17x-13?

OpenStudy (anonymous):

That's a monster!

OpenStudy (anonymous):

First check if \(2\) is a root.

OpenStudy (anonymous):

Check \(1, -1, 2, -2\) and pray one of them is a root.

OpenStudy (anonymous):

If one of them is a root then divide out \(x-\text{root}\)

OpenStudy (wolfe8):

I think you're supposed to use factorization.

OpenStudy (anonymous):

If you see how to do that @wolfgirl then show how.

OpenStudy (anonymous):

i did use syntethetic division and use -1 and 1/2 as the divisor then after that i get the answer 2x^2-8x+26 then since i couldny use factorixation then i use quadratic formula and my answers turn out to be 10sqrtof11all over 4 and 6sqrt of 11 all over 4

OpenStudy (anonymous):

and the problem is idki if thats was correct ill appreciate if youll tell me my errors

OpenStudy (anonymous):

Show your work

OpenStudy (wolfe8):

Square both sides: 2 x^4-7 x^3+21 x^2+17 x-13 = 0 The left hand side factors into a product with three terms: (x+1) (2 x-1) (x^2-4 x+13) = 0 I think you can do the rest. Btw, I'm not Wolfgirl :)

OpenStudy (wolfe8):

Also, you made a mistake in the question with that #. I know from chat though you meant 3

OpenStudy (tkhunny):

Don't check 2 or -2. The Rational Root Theorem suggests +/- 1, +/- 1/2, +/- 13, and +/- 13/2 By your result, you had hits on x = -1 and x = 1/2.

OpenStudy (anonymous):

@tkhunny so is my results correct?

OpenStudy (tkhunny):

Typically, I would prefer (2x-1) in the place of (x-1/2), but other than that it seems fine.